OpenStudy (anonymous):

Perfect Square Trinomials: Factor Completely : 144r^2-24r+1

OpenStudy (anonymous):

(12r-1)^2

OpenStudy (anonymous):

Thanks, I was confused. But I understand now.

OpenStudy (anonymous):

144=12*12 1=1*1 and u just gotta take the negative sign into account (12r-1)^2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):